OpenStudy (anonymous):

Arrange the fractions in ascending order. Or, if the fractions are equivalent, answer "Equivalent". ,

OpenStudy (anonymous):

1/5 & 13/17

OpenStudy (anonymous):

\[\frac{1}{5}<\frac{13}{17}\]

OpenStudy (saifoo.khan):

\[\checkmark\frac{1}{5}<\frac{13}{17}\checkmark\]

OpenStudy (anonymous):

thanks